# Robust Beamforming Design and Antenna Selection for Dynamic HRIS-aided Massive MIMO Systems

### v1

@@ Abstract
In this paper, a dynamic hybrid active-passive reconfigurable intelligent surface (HRIS) is proposed to further enhance the massive multiple-input-multiple-output (MIMO) system, since it supports the dynamic placement of active and passive elements. Specifically, considering the impact of the hardware impairments (HWIs), we investigate the channel-aware configuration of the receive antennas at the base station (BS) and the active/passive elements at the HRIS to improve the reliability of system. To this end, we investigate the average mean-square-error (MSE) minimization problem for the HRIS-aided massive MIMO system by jointly optimizing the BS receive antenna selection matrix, the reflection phase coefficients, the reflection amplitude matrix, and the mode selection matrix of the HRIS under the power budget of the HRIS. To tackle the non-convexity and intractability of this problem, we first transform the binary and discrete variables into continuous ones, and then propose a penalty-based exact block coordinate descent (BCD) algorithm to solve these subproblems alternately. Numerical simulations demonstrate the great superiority of the proposed scheme over the conventional benchmark schemes.
@@ I Introduction
[lines=2]MASSIVE multiple-input multiple-output (MIMO) has been a revolutionary solution for enhancing the spectral efficiency (SE) to meet the rapidly growing traffic demand for future wireless communication systems. However, fully digital implementation of massive MIMO requires costly dedicated radio-frequency (RF) chains for each antenna, thus inevitably resulting in prohibitive hardware cost and power consumption.
@@
To improve both cost and power efficiency, analog signal processing can be additionally introduced to reduce the number of RF chains [1]. One common technique is the so-called hybrid analog-digital transceiver widely used in recent massive MIMO systems. The signal undergoes initial processing by a low-dimensional digital beamformer before traversing through a high-dimensional analog beamformer, constructed using analog phase shifters. Generally, there are two typical structure for implementing the analog signal processing. By connecting the RF chains to all antennas or a subset of the antenna array, the fully-connected or sub-connected hybrid analog-digital transceiver is utilized [2]. However, hybrid beamforming still requires a large number of phase shifters, whereas high-resolution ones are generally associated with complicated circuits and high energy consumption in practice. Another low-cost alternative technique is antenna selection [3, 4], where only a subarray of antennas is activated with the limited RF chains via the RF switch network. Compared to the phase shifters, the analog RF switches have a much lower complexity and power consumption, making the antenna selection more appealing [5].
On a parallel track, reconfigurable intelligent surface (RIS) has been envisioned as a cost-effective solution for realizing green communication in 6G [6]. RIS exhibits the remarkable capability to dynamically reshape the wireless propagation environment and explore new physical dimensions of transmission. By deploying a massive number of low-cost passive elements, RIS can dynamically adjust the phase of the incident signal to create favorable wireless channels, thereby improving the communication performance. However, due to the fully passive nature of RIS elements, the widespread adoption of RIS has been greatly constrained. For example, the performance improvement induced by RIS is limited due to the “multiplicative fading” effect, particularly in the presence of a direct link. To overcome this challenge, the hybrid active-passive RIS (HRIS) has been proposed to compensate for the severe cascaded path loss with additional power amplifiers (PAs) [7, 8]. In [7], the authors investigated the optimal ratio between the number of active and passive elements to maximize the ergodic capacity under the total power budget. On the other hand, the authors in [8] have studied the channel-aware placement of the active elements to enhance the signal-to-noise ratio (SNR). Unfortunately, these existing works focused on the ideal hardware implementation and ignored the impact of hardware impairments (HWIs).
@@
In this paper, we investigate a dynamic HRIS-assisted massive MIMO system under the non-ideal hardware implementation. The channel-aware placement of the base station (BS) receive antennas and the active/passive HRIS elements are optimized to enhance the system performance. To this end, we investigate the average mean-square-error (MSE) minimization problem under the power budget of HRIS by jointly optimizing the receive antenna selection matrix, the reflection phase coefficients, the reflection amplitude matrix, and the mode selection matrix for the active and passive elements. To tackle the mix-integer optimization problem, a penalty-based exact block coordinate descent (BCD) algorithm is proposed. Numerical simulations show great superiority in the channel-aware configuration of the BS receive antennas and the active/passive HRIS elements as compared to the conventional benchmarks.
@@ II-A System Model
As depicted in Fig. 1, a dynamic HRIS-aided uplink multi-user communication system with $M$ single-antenna users and an $N_{R}$-antenna BS is considered. The time division multiple access (TDMA) scheme is adopted, where the users are served by the BS over orthogonal time slots. The HRIS consists of $N$ elements, where the active elements are equipped with power amplifiers. To reduce the power consumption and hardware complexity, we assume the BS receives data streams from $N_{R}$ antennas with $L<N_{R}$ dedicated RF chains. The antenna selection technique is employed to select $L$ out of $N_{R}$ antennas to explore the spatial diversity.
@@
We denote ${\bf{h}}_{d}\in\mathbb{C}^{N_{R}\times 1}$ and ${\bf{G}}\in\mathbb{C}^{N\times N_{R}}$ as the full channel state information (CSI) associated with $N_{R}$ receive antennas from the BS to the user and the HRIS, respectively. The channel between the HRIS and the user is ${\bf{h}}_{r}\!\!\in\!\!\mathbb{C}^{N\times 1}$. Based on the receive antenna selection, $L$ out of $N_{R}$ or $N$ rows are selected to form the ${L\times 1}$ or ${L\times N}$ sub-channels, respectively. Denoting ${\bf{A}}\in\mathcal{A}$ as the index matrix for the $L$ selected antennas, where $\mathcal{A}\triangleq\{{\bf{A}}\in\mathbb{C}^{L\times N_{R}};A_{i,j}\in\{0,1\};\sum\nolimits_{j=1}^{N_{R}}A_{i,j}\!=\!1,\forall i;\sum\nolimits_{i=1}^{L}A_{i,j}\!\leq\!1,\forall j\}$, the sub-channels can be represented by ${\bf{A}}{\bf{h}}_{d}$ and ${\bf{A}}{\bf{G}}$. All channels are assumed to experience quasi-static flat fading and perfectly available using the recent advances in RIS and massive MIMO channel estimation. [^1] Due to the spatial correlation across the large-scale antenna array, the full-dimensional CSI can be recovered from the partial CSI obtained via pilot training [9].
Figure 1: A dynamic HRIS-assisted multi-user uplink massive MIMO communication system.
@@
The HRIS comprises $N_{a}$ active and $N_{p}$ passive elements, where $N_{a}+N_{p}=N$. The location and the number of active elements are optimized to enhance the system performance. Specifically, we denote ${\bm{\Phi}}\triangleq{\rm diag}(e^{j\phi_{1}},\dots,e^{j\phi_{N}})$ as the reflection phase matrix, where $e^{j\phi_{n}}$ denotes the discrete phase shift of the $n$-th element with $\phi_{n}\in\mathcal{X}\triangleq\{0,\frac{2\pi}{2^{B}},\dots,\frac{2\pi(2^{B}-1)}{2^{B}}\}$. $B$ is the number of quantization bits. The reflection amplitude matrix of HRIS is denoted as ${\bf{\Lambda}}\triangleq{\rm diag}(\omega_{1},\dots,\omega_{N})$, where $\omega_{n}$ represents the amplification factor of $n$-th element. For the active elements, we have $\omega_{n}>{\mu}_{\rm min}$ when $n\in\mathcal{N}_{a}$, where ${\mu}_{\rm min}\geq 1$ denotes the minimum amplification factor and $\mathcal{N}_{a}$ is the index set of all active elements. Otherwise, we have $\omega_{n}=1$ for those passive elements. The mode selection matrix related to $\mathcal{N}_{a}$ is denoted as ${\bf{B}}={\rm diag}({\bm{\gamma}})\in\mathcal{B}\triangleq\{{\bm{\gamma}}\in\mathbb{C}^{N\times 1};\gamma_{i}=1,\forall i\in\mathcal{N}_{a};\gamma_{i}=0,\forall i\notin\mathcal{N}_{a}\}$. For simplicity, we assume all the active elements share one common amplification factor $\mu$.
@@
With the non-ideal phase shifters of HRIS, the phase error of the $n$-th element can be modeled as ${\bar{\phi}}_{n}\in\mathcal{U}[-\frac{\pi}{2^{B}},\frac{\pi}{2^{B}}]$, where $\mathcal{U}[\cdot]$ denotes the uniform distribution [10]. Therefore, the actual phase shift matrix with phase noise is ${\bm{\tilde{\Phi}}}={{\bm{\Phi}}}{\bm{\bar{\Phi}}}$, with ${\bm{\bar{\Phi}}}\triangleq{\rm diag}(e^{j{\bar{\phi}}_{1}},\dots,e^{j{\bar{\phi}}_{N}})$ denoting the phase noise matrix.
To accurately model the realistic imperfect hardware, the additive HWIs are considered. Specifically, the aggregate residual HWIs for the transceiver can be modeled as the additive Gaussian distortion noise [11]. As such, the actual transmitted signal for the user is expressed as
@@
where ${{\bf{y}}}$ denotes the un-distorted received signal and ${\bf{n}}_{b}$ stands for the additive white Gaussian noise (AWGN) whose elements are random variables with zero-mean and unit-variance, i.e., ${\bf{n}}_{b}\!\sim\!\mathcal{CN}({\bf{0}},{\sigma_{b}^{2}}{\bf{I}}_{L})$. ${\bf{G}}^{H}{\bf{B}}{\bf{\Lambda}}{\bm{\tilde{\Phi}}}{\bf{n}}_{a}$ in ${{\bf{y}}}$ represents the amplified noise associated with the active elements, where ${\bf{n}}_{a}\!\sim\!\mathcal{CN}({\bf{0}},{\sigma_{a}^{2}}{\bf{I}}_{N})$. ${\bm{\kappa}}_{r}$ stands for the receive distortion noise. We have ${\bm{\kappa}}_{r}\!\sim\!\mathcal{CN}(({\bf{0}},{k_{r}^{2}}{\rm\widetilde{diag}}(\mathbb{E}[{\bf{y}}{\bf{y}}^{H}]))$, whose variance is proportional to the average received power of the individual antennas at the BS. $k_{r}$ represents the received distortion level and $\mathbb{E}[{\bf{y}}{\bf{y}}^{H}]$ is derived as
@@
Assume that the receive beamforming at the BS is ${\bf{w}}$, then the estimated symbol is denoted as
@@
with ${\bf{\tilde{h}}}={\bf{h}}_{d}+{\bf{G}}^{H}{\bf{\Lambda}}{\bm{\tilde{\Phi}}}{\bf{h}}_{r}$. Therefore, the average MSE can be derived as
@@ II-B Problem formulation
In this paper, we mainly focus on enhancing the transmission reliability of the considered system in terms of MSE. Specifically, the average MSE is minimized under the total power budget $P_{\rm HRIS}$ of HRIS by jointly optimizing the receive antenna selection matrix ${\bf{A}}$, the mode selection matrix ${\bf{B}}$, the reflection phase matrix ${\bf{\Phi}}$, and the reflection amplitude matrix ${\bf{\Lambda}}$. The total power consumption at the HRIS is calculated as $P\!=\!\mathbb{E}[|{\bf{\Lambda}}{\bf{B}}{\bm{\tilde{\Phi}}}{\bf{h}}_{r}{\tilde{s}}|^{2}]\!+\!\mathbb{E}[|{\bf{B}}{\bf{\Lambda}}{\bm{\tilde{\Phi}}}{\bf{n}}_{a}|^{2}]={\tilde{p}}|{\bf{\Lambda}}{\bf{B}}{\bf{h}}_{r}|^{2}+\sigma_{a}^{2}|{\bf{B}}{\bf{\Lambda}}|^{2}$. Then, the average MSE minimization problem can be formulated as
@@
$\quad\qquad\phi_{n}\in\mathcal{X},$ (6c)
$\quad\qquad\omega_{n}>\mu_{\rm min},\forall n\in\mathcal{N}_{a},$ (6d)
$\quad\qquad{\bf{A}}\in\mathcal{A},{\bf{B}}\in\mathcal{B},\par$ (6e)
where (6c) denotes the discrete phase coefficient constraint and (6d) limits the amplification factor of the active elements. Unfortunately, the problem (P1) is intractable and NP-hard due to the integer constraints and tightly coupled variables. In the sequel, we first transform these intractable constraints and then present a penalty-based exact BCD algorithm to obtain a stationary solution.
III Proposed exact BCD algorithm

### v2

# Robust Beamforming Design and Antenna Selection for Dynamic HRIS-aided MISO Systems

@@ Abstract
In this paper, we propose a dynamic hybrid active-passive reconfigurable intelligent surface (HRIS) to enhance multiple-input-single-output (MISO) communications, leveraging the property of dynamically placing active elements. Specifically, considering the impact of hardware impairments (HWIs), we investigate channel-aware configurations of the receive antennas at the base station (BS) and the active/passive elements at the HRIS to improve transmission reliability. To this end, we address the average mean-square-error (MSE) minimization problem for the HRIS-aided MISO system by jointly optimizing the BS receive antenna selection matrix, the reflection phase coefficients, the reflection amplitude matrix, and the mode selection matrix of the HRIS. To overcome the non-convexity and intractability of this problem, we first transform the binary and discrete variables into continuous ones, and then propose a penalty-based exact block coordinate descent (PEBCD) algorithm to alternately solve these subproblems. Numerical simulations demonstrate the significant superiority of our proposed scheme over conventional benchmark schemes.
@@ I Introduction
[lines=2]MASSIVE multiple-input multiple-output (MIMO) has been a revolutionary solution for enhancing spectral efficiency (SE) to meet the rapidly growing traffic demand for future wireless communication systems. However, a fully digital implementation of MIMO requires costly dedicated radio-frequency (RF) chains for each antenna, inevitably resulting in prohibitive hardware costs and power consumption.
@@
To improve both cost and power efficiency, analog signal processing can be additionally introduced to reduce the number of RF chains [1]. One common technique is the so-called hybrid analog-digital transceiver, widely used in recent massive MIMO systems. The signal undergoes initial processing by a low-dimensional digital beamformer before traversing through a high-dimensional analog beamformer, constructed using analog phase shifters. Generally, there are two typical structures for implementing analog signal processing. By connecting the RF chains to all antennas or a subset of the antenna array, fully-connected or sub-connected hybrid analog-digital transceivers are utilized [2]. However, hybrid beamforming still requires a large number of phase shifters, which are generally associated with complicated circuits and high energy consumption in practice. Another low-cost alternative technique is antenna selection [3], where only a subarray of antennas is activated with the limited RF chains via the RF switch network. Compared to phase shifters, analog RF switches have much lower complexity and power consumption, making antenna selection more appealing [4].
On a parallel track, advances in metasurfaces have promoted a new paradigm in wireless communications [5]. For example, authors in [6] propose stacked intelligent metasurfaces (SIM) at the transmitter and receiver to implement holographic multiple-input multiple-output (HMIMO) communications without requiring excessive RF chains. Although metasurface-aided transceivers have huge potential for high energy efficiency, they still face challenges in the absence of line-of-sight scenarios. On the other hand, reconfigurable intelligent surface (RIS) [7] has exhibited remarkable capability to dynamically reshape the wireless propagation environment and explore new physical dimensions of transmission. By deploying a massive number of low-cost passive elements, RIS can dynamically adjust the phase of the incident signal to create favorable wireless channels, thereby improving communication performance. However, due to the fully passive nature of RIS elements, the widespread adoption of RIS has been greatly constrained. For example, the performance improvement induced by RIS is limited due to the ”multiplicative fading” effect, particularly in the presence of a direct link.
To overcome this challenge, the hybrid active-passive RIS (HRIS) has been proposed to compensate for the severe cascaded path loss with additional power amplifiers (PAs) [8, 9]. Unlike the amplify-and-forward (AF) relay, HRIS operates similarly to conventional RIS by directly reflecting the incident signal with the desired adjustments at the electromagnetic (EM) level. In contrast, an AF relay typically requires larger and more power-hungry RF chains to receive and then transmit the signal with amplification. This process usually occurs at the baseband level and requires two time slots to complete the amplify-and-forward cycle. Even when operating in full-duplex mode, the AF relay increases hardware complexity due to the need to mitigate self-interference. Additionally, the AF relay simply amplifies the received signal without modifying the phase. In [8], the authors investigated the optimal ratio between the number of active and passive elements to maximize the ergodic capacity under the total power budget. On the other hand, the authors in [9] have studied the channel-aware placement of the active elements to enhance the signal-to-noise ratio (SNR). Unfortunately, these existing works focused on the ideal hardware implementation and ignored the impact of hardware impairments (HWIs).
@@
In this paper, we investigate a dynamic HRIS-assisted MISO communication system under the non-ideal hardware implementationThe channel-aware placement of the base station (BS) receive antennas and the active/passive HRIS elements are optimized to enhance the system performance. To this end, we investigate the average mean-square-error (MSE) minimization problem under the power budget of HRIS by jointly optimizing the receive antenna selection matrix, the reflection phase coefficients, the reflection amplitude matrix, and the mode selection matrix for the active and passive elements. To tackle the mix-integer optimization problem, a penalty-based exact block coordinate descent (PEBCD) algorithm is proposed. Numerical simulations show great superiority in the channel-aware configuration of the BS receive antennas and the active/passive HRIS elements as compared to the conventional benchmarks.
Figure 1: A dynamic HRIS-assisted uplink MISO communication system.
@@ II-A System Model
As depicted in Fig. 1, a dynamic HRIS-aided uplink MISO communication system with a single-antenna user and an $N_{R}$-antenna BS is considered. The HRIS comprises $N_{a}$ active and $N_{p}$ passive elements, where $N_{a}+N_{p}=N$. To reduce the power consumption and hardware complexity, we assume the BS receives the data stream from $N_{R}$ antennas with $L<N_{R}$ dedicated RF chains. The antenna selection technique is employed to select $L$ out of $N_{R}$ antennas to explore the spatial diversity.
@@
We denote ${\bf{h}}_{d}\in\mathbb{C}^{N_{R}\times 1}$ and ${\bf{G}}\in\mathbb{C}^{N\times N_{R}}$ as the full channel state information (CSI) associated with $N_{R}$ receive antennas from the BS to the user and the HRIS, respectively. The channel between the HRIS and the user is ${\bf{h}}_{r}\!\!\in\!\!\mathbb{C}^{N\times 1}$. Based on the receive antenna selection, $L$ out of $N_{R}$ or $N$ rows are selected to form the ${L\times 1}$ or ${L\times N}$ sub-channels, respectively. Denoting ${\bf{A}}\in\mathcal{A}$ as the index matrix for the $L$ selected antennas, where $\mathcal{A}\triangleq\{{\bf{A}}\in\mathbb{C}^{L\times N_{R}};A_{i,j}\in\{0,1\};\sum\nolimits_{j=1}^{N_{R}}A_{i,j}\!=\!1,\forall i;\sum\nolimits_{i=1}^{L}A_{i,j}\!\leq\!1,\forall j\}$, the sub-channels can be represented by ${\bf{A}}{\bf{h}}_{d}$ and ${\bf{A}}{\bf{G}}$. All channels are assumed to experience quasi-static flat fading and perfectly available using the recent advances in RIS and MISO channel estimation. [^1] To delve into the unique effects of hardware impairments on the dynamic HRIS-assisted MISO communication system, we adopt the ideal CSI assumption, which is prevalent in contemporary studies [10, 11]. The development of a channel estimation method falls outside the scope of this work and is therefore reserved for future exploration. It is crucial to highlight that the optimization framework we propose is designed to be easily adaptable to scenarios characterized by imperfect CSI.
@@
The dynamic HRIS, as shown in Fig. 1, is equipped with low-cost discrete phase shifters and RF switches. To avoid the bulky circuits in practical implementation, we adopt the “sub-connected architecture” for those active HRIS elements. This approach differs from the fully-connected one by employing a single power amplifier to serve all elements. The RF switches are used in conjunction with different phase-shift circuits, allowing HRIS elements to operate in either active or passive modes and independently control phase shifts while sharing a common amplification factor. The location and the number of active elements are optimized to enhance the system performance. Specifically, we denote ${\bm{\Phi}}\triangleq{\rm diag}(e^{j\phi_{1}},\dots,e^{j\phi_{N}})$ as the reflection phase matrix, where $e^{j\phi_{n}}$ denotes the discrete phase shift of the $n$-th element with $\phi_{n}\in\mathcal{X}\triangleq\{0,\frac{2\pi}{2^{B}},\dots,\frac{2\pi(2^{B}-1)}{2^{B}}\}$. $B$ is the number of quantization bits. The reflection amplitude matrix of HRIS is denoted as ${\bf{\Lambda}}\triangleq{\rm diag}(\omega_{1},\dots,\omega_{N})$, where $\omega_{n}$ represents the amplification factor of $n$-th element. For the active elements, we have $\omega_{n}>{\mu}_{\rm min}$ when $n\in\mathcal{N}_{a}$, where ${\mu}_{\rm min}\geq 1$ denotes the minimum amplification factor and $\mathcal{N}_{a}$ is the index set of all active elements. Otherwise, we have $\omega_{n}=1$ for those passive elements. With the same PA, all the active elements share one common amplification factor $\mu$. We introduce a binary diagonal selection matrix ${\bf{B}}$ to establish the relationship between the HRIS elements and the PA, i.e., ${\bf{B}}={\rm diag}({\bm{\gamma}})\in\mathcal{B}\triangleq\{{\bm{\gamma}}\in\mathbb{C}^{N\times 1};\gamma_{i}=1,\forall i\in\mathcal{N}_{a};\gamma_{i}=0,\forall i\notin\mathcal{N}_{a}\}$.
@@
Due to the non-ideal phase shifters of HRIS, the phase error of the $n$-th element can be modeled as ${\bar{\phi}}_{n}\in\mathcal{U}[-\frac{\pi}{2^{B}},\frac{\pi}{2^{B}}]$, where $\mathcal{U}[\cdot]$ denotes the uniform distribution [12]. Therefore, the actual phase shift matrix with phase noise is ${\bm{\tilde{\Phi}}}={{\bm{\Phi}}}{\bm{\bar{\Phi}}}$, with ${\bm{\bar{\Phi}}}\triangleq{\rm diag}(e^{j{\bar{\phi}}_{1}},\dots,e^{j{\bar{\phi}}_{N}})$ denoting the phase noise matrix. To accurately model the realistic imperfect hardware, the additive HWIs are considered. Specifically, the aggregate residual HWIs for the transceiver can be modeled as the additive Gaussian distortion noise [13]. As such, the actual transmitted signal for the user is expressed as
@@
where ${{\bf{y}}}$ denotes the un-distorted received signal and ${\bf{n}}_{b}$ stands for the additive white Gaussian noise (AWGN) whose elements are random variables with zero-mean and unit-variance, i.e., ${\bf{n}}_{b}\!\sim\!\mathcal{CN}({\bf{0}},{\sigma_{b}^{2}}{\bf{I}}_{L})$. ${\bf{G}}^{H}{\bf{B}}{\bf{\Lambda}}{\bm{\tilde{\Phi}}}{\bf{n}}_{a}$ in ${{\bf{y}}}$ represents the amplified noise associated with the active elements, where ${\bf{n}}_{a}\!\sim\!\mathcal{CN}({\bf{0}},{\sigma_{a}^{2}}{\bf{I}}_{N})$. ${\bm{\kappa}}_{r}$ stands for the receive distortion noise. We have ${\bm{\kappa}}_{r}\!\sim\!\mathcal{CN}(({\bf{0}},{k_{r}^{2}}{\rm\widetilde{diag}}(\mathbb{E}[{\bf{y}}{\bf{y}}^{H}]))$, whose variance is proportional to the average received power of the individual antennas at the BS. ${\rm diag}(\cdot)$ returns a square diagonal matrix whose diagonal elements are the same as the input. $k_{r}$ represents the received distortion level and $\mathbb{E}[{\bf{y}}{\bf{y}}^{H}]$ is derived as
@@
Assume that the receive beamforming at the BS is ${\bf{w}}$, then the estimated symbol can be denoted as
@@
with ${\bf{\tilde{h}}}={\bf{h}}_{d}+{\bf{G}}^{H}{\bf{\Lambda}}{\bm{\tilde{\Phi}}}{\bf{h}}_{r}$. Therefore, the average MSE222 Due to the correlation between the minimum MSE and the minimum average bit error rate (BER), the MSE metric provide a elegant expression for measuring the quality of service compared to the BER, which is difficult or intractable due to their dependence on the integral $Q$ function [14]. can be derived as
@@ II-B Problem formulation
In this paper, we mainly focus on enhancing transmission reliability of the considered system in terms of MSE. Specifically, the average MSE is minimized under the total power budget $P_{\rm HRIS}$ of HRIS by jointly optimizing the receive antenna selection matrix ${\bf{A}}$, the mode selection matrix ${\bf{B}}$, the reflection phase matrix ${\bf{\Phi}}$, and the reflection amplitude matrix ${\bf{\Lambda}}$. The total power consumption at the HRIS is calculated as $P\!=\!\mathbb{E}[|{\bf{\Lambda}}{\bf{B}}{\bm{\tilde{\Phi}}}{\bf{h}}_{r}{\tilde{s}}|^{2}]\!+\!\mathbb{E}[|{\bf{B}}{\bf{\Lambda}}{\bm{\tilde{\Phi}}}{\bf{n}}_{a}|^{2}]={\tilde{p}}|{\bf{\Lambda}}{\bf{B}}{\bf{h}}_{r}|^{2}+\sigma_{a}^{2}|{\bf{B}}{\bf{\Lambda}}|^{2}$. Then, the average MSE minimization problem can be formulated as
@@
$\quad\qquad\qquad\phi_{n}\in\mathcal{X},$ (6c)
$\quad\qquad\qquad\omega_{n}>\mu_{\rm min},\forall n\in\mathcal{N}_{a},$ (6d)
$\quad\qquad\qquad{\bf{A}}\in\mathcal{A},{\bf{B}}\in\mathcal{B},\par$ (6e)
where (6c) denotes the discrete phase coefficient constraint and (6d) limits the amplification factor of the active elements. Unfortunately, the problem (P1) is intractable and NP-hard due to the integer constraints and tightly coupled variables. In the sequel, we first transform these intractable constraints and then present a PEBCD algorithm to obtain a stationary solution.
III Proposed PEBCD algorithm